A mixture of air and water vapour at 1.013 bar and 16 degree C has a dew
point of 5degree C. Calculate the relative and specific humidity’s?
1
Expert's answer
2018-12-17T07:24:23-0500

The saturaion pressures at the given temperatures are

"P_s(5^o C)=872.6 Pa,\\
P_s(16^o C)=1818.8 Pa,"

The relative humudity is given by

"\phi=\frac{P_s(5^o C)}{P_s(16^o C)}\\
\phi= \frac{872.6}{1818.8}=0.480"

The specific humidity is given by

"x=\frac{P_s(5^o C)}{P-P_s(5^o C)}\\
x= \frac{872.6}{1.013\times10^5-872.6}=0.0054 kg/kg."